OpenStudy (anonymous):

For 2x - 3y = -10, when x = 4. What is the value of y?

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

2x-3y=-10 2(4)-3y=-10 8-3y=-10 add 10 and add 3y 18=3y divide by 3 both sides y=5